OpenStudy (anonymous):

solve cosθ + 3sin(θ/2) -2 =o for 0≤θ≤π

OpenStudy (freckles):

you can use the double angle identity for cosine

OpenStudy (freckles):

\[\cos(\theta)=\cos^2(\frac{\theta}{2})-\sin^2(\frac{\theta}{2}) \\ \text{ and by pythagorean identity } \\ \cos(\theta)=1-2 \sin^2(\frac{\theta}{2})\]

OpenStudy (freckles):

using this as a substitution will give you a quadratic in terms of sin(theta/2)

OpenStudy (anonymous):

thanks :))

OpenStudy (freckles):

np let me know if you need further help